# Fairness in Large Language Models: A Taxonomic Survey

### v1

@@ 6.2. In-training
i) Loss Function Modification. Loss function modification involves incorporating a fairness-constrained into the training process of downstream tasks to guide the model toward fair learning. Wang et al. (Wang et al., 2021a) introduced an approach that integrates causal relationships into model training. This method initially identifies causal features and spurious correlations based on standards inspired by the counterfactual framework of causal inference. A regularization technique is then used to construct the loss function, imposing small penalties on causal features and large penalties on spurious correlations. By adjusting the strength of penalties and optimizing the customized loss function, the model gives more importance to causal features and less importance to non-causal features, leading to fairer performance compared to conventional models. Additionally, Park et al. (Park et al., 2023) proposed an embedding-based objective function that addresses the persistence of gender-related features in stereotype word vectors by utilizing generated gender direction vectors during fine-tuning steps.
ii) Auxiliary Module. Auxiliary module involve the addition of modules with the purpose of reducing bias within the model structure to help diminish bias. For instance, Lauscher et al. (Lauscher et al., 2021) proposed a sustainable modular debiasing strategy, namely Adapter-based DEbiasing of LanguagE Models (ADELE). Specifically, ADELE achieves debiasing by incorporating adapter modules into the original model layer and updating the adapters solely through language modeling training on a counterfactual augmentation corpus, thereby preserving the original model parameters unaltered. Additionally, Shen et al. (Ravfogel et al., 2020) introduces Iterative Null Space Projection (INLP) for removing information from neural representations. Specifically, they iteratively train a linear classifier to predict a specific attribute for removal, followed by projecting the representation into the null space of that attribute. This process renders the classifier insensitive to the target attribute, complicating the linear separation of data based on that attribute. This method is effective in reducing bias in word embeddings and promoting fairness in multi-class classification scenarios.
@@ 6.4. Post-processing
Post-processing approaches modify the results generated by the model to mitigate biases, which is particularly crucial for closed-source LLMs where obtaining probabilities and embeddings of generated text is challenging, limiting the direct modification to output results only. Here, the method of chain-of-thought and rewriting serve as the illustrative approaches to convey this concept.
i) Chain-of-thought (CoT). The CoT technique enhances the hopeful and performance of LLMs towards fairness by leading them through incremental reasoning steps. The work by Kaneko et al. (Kaneko et al., 2024) provided a benchmark test where LLMs were tasked with determining the gender associated with specific occupational terms. Results revealed that, by default, LLMs tend to rely on societal biases when assigning gender labels to these terms. However, incorporating CoT prompts mitigates these biases. Furthermore, Dhingra et al. (Dhingra et al., 2023) introduced a technique combining CoT prompts and SHAP analysis (Lundberg and Lee, 2017) to counter stereotypical language towards queer individuals in model outputs. Using SHAP, stereotypical terms related to LGBTQ+ [^6] https://en.wikipedia.org/wiki/LGBT individuals were identified, and then the chain-of-thought approach was used to guide language models in correcting this language.
@@ 7.2. Datasets
i) Probability-based. As mentioned in section 5.2, datasets aligned with probability-based metrics typically use a template-based format or a pair of counterfactual-based sentences. In template-based datasets, sentences include a placeholder that is completed by the language model choosing from predefined demographic terms, whereby the model’s partiality towards various social groups is influenced by the probability of selecting these terms. Noteworthy examples of such datasets include WinoBias (Zhao et al., 2018), which assess a model’s competence in linking gender pronouns and occupations in both stereotypical and counter-stereotypical scenarios. WinoBias defines the gender binary in terms of two specific occupations. Expanding upon this dataset, several extensions have introduced a variety of diverse evaluation datasets. For example, WinoBias+ (Vanmassenhove et al., 2021) enhances the original WinoBias dataset by employing rule-based and neural neutral rewriters to convert gendered sentences into neutral equivalents. Additionally, BUG (Levy et al., 2021) broadens the evaluation of gender bias in machine translation by using a large-scale real-world English dataset. In contrast, GAP (Webster et al., 2018) introduces a gender-balanced tagged corpus comprising 8,908 ambiguous pronoun-name pairs, providing a more balanced dataset for accurately assessing model bias. Another category of counterfactual-based datasets evaluates bias by presenting the model with pairs of sentences containing different demographic terms and assessing their likelihood using a pseudo-log-likelihood metric. For instance, CrowS-Pairs (Nangia et al., 2020) provides 1,508 sentence pairs to assess stereotypes of historically marginalized social groups, addressing nine biases related to race, gender, sexual orientation, religion, age, nationality, disability, physical appearance, and socioeconomic status. Similarly, WinoQueer (Felkner et al., 2023), developed by the LGBTQ+ community, contains 45,540 sentence pairs designed to gauge anti-LGBTQ+ stereotypes, encompassing LGBTQ+ identity descriptors and counterfactual versions without LGBTQ+ identity markers.
@@ 9. Conclusion
LLMs have demonstrated remarkable success across various high-impact applications, transforming the way we interact with technology. However, without proper fairness safeguards, they risk making decisions that could lead to discrimination, presenting a serious ethical issues and an increasing societal concern. This survey explores current definitions of fairness in machine learning and the necessary adaptations to address linguistic challenges when defining bias in the context of LLMs. Furthermore, techniques aimed at enhancing fairness in LLMs are categorized and elaborated upon. Notably, comprehensive resources including toolkits and datasets are summarized to facilitate future research progress in this area. Finally, existing challenges and open questions areas are also discussed.

### v2

@@ 6.2. In-training
i) Loss Function Modification. Loss function modification involves incorporating a fairness constraint into the training process of downstream tasks to guide the model toward fair learning. Wang et al. (Wang et al., 2021a) introduced an approach that integrates causal relationships into model training. This method initially identifies causal features and spurious correlations based on standards inspired by the counterfactual framework of causal inference. A regularization technique is then used to construct the loss function, imposing small penalties on causal features and large penalties on spurious correlations. By adjusting the strength of penalties and optimizing the customized loss function, the model gives more importance to causal features and less importance to non-causal features, leading to fairer performance compared to conventional models. Additionally, Park et al. (Park et al., 2023) proposed an embedding-based objective function that addresses the persistence of gender-related features in stereotype word vectors by utilizing generated gender direction vectors during fine-tuning steps.
ii) Auxiliary Module. Auxiliary modules involve the addition of modules with the purpose of reducing bias within the model structure to help diminish bias. For instance, Lauscher et al. (Lauscher et al., 2021) proposed a sustainable modular debiasing strategy, namely Adapter-based DEbiasing of LanguagE Models (ADELE). Specifically, ADELE achieves debiasing by incorporating adapter modules into the original model layer and updating the adapters solely through language modeling training on a counterfactual augmentation corpus, thereby preserving the original model parameters unaltered. Additionally, Shen et al. (Ravfogel et al., 2020) introduces Iterative Null Space Projection (INLP) for removing information from neural representations. Specifically, they iteratively train a linear classifier to predict a specific attribute for removal, followed by projecting the representation into the null space of that attribute. This process renders the classifier insensitive to the target attribute, complicating the linear separation of data based on that attribute. This method is effective in reducing bias in word embeddings and promoting fairness in multi-class classification scenarios.
@@ 6.4. Post-processing
Post-processing approaches modify the results generated by the model to mitigate biases, which is particularly crucial for closed-source LLMs where obtaining probabilities and embeddings of generated text is challenging, limiting the direct modification to output results only. Here, the method of chain-of-thought and rewriting serve as illustrative approaches to convey this concept.
i) Chain-of-thought (CoT). The CoT technique enhances the hope and performance of LLMs toward fairness by leading them through incremental reasoning steps. The work by Kaneko et al. (Kaneko et al., 2024) provided a benchmark test where LLMs were tasked with determining the gender associated with specific occupational terms. Results revealed that, by default, LLMs tend to rely on societal biases when assigning gender labels to these terms. However, incorporating CoT prompts mitigates these biases. Furthermore, Dhingra et al. (Dhingra et al., 2023) introduced a technique combining CoT prompts and SHAP analysis (Lundberg and Lee, 2017) to counter stereotypical language towards queer individuals in model outputs. Using SHAP, stereotypical terms related to LGBTQ+ [^6] https://en.wikipedia.org/wiki/LGBT individuals were identified, and then the chain-of-thought approach was used to guide language models in correcting this language.
@@ 7.2. Datasets
i) Probability-based. As mentioned in section 5.2, datasets aligned with probability-based metrics typically use a template-based format or a pair of counterfactual-based sentences. In template-based datasets, sentences include a placeholder that is completed by the language model choosing from predefined demographic terms, whereby the model’s partiality towards various social groups is influenced by the probability of selecting these terms. Noteworthy examples of such datasets include WinoBias (Zhao et al., 2018), which assesses a model’s competence in linking gender pronouns and occupations in both stereotypical and counter-stereotypical scenarios. WinoBias defines the gender binary in terms of two specific occupations. Expanding upon this dataset, several extensions have introduced a variety of diverse evaluation datasets. For example, WinoBias+ (Vanmassenhove et al., 2021) enhances the original WinoBias dataset by employing rule-based and neural-neutral rewriters to convert gendered sentences into neutral equivalents. Additionally, BUG (Levy et al., 2021) broadens the evaluation of gender bias in machine translation by using a large-scale real-world English dataset. In contrast, GAP (Webster et al., 2018) introduces a gender-balanced tagged corpus comprising 8,908 ambiguous pronoun-name pairs, providing a more balanced dataset for accurately assessing model bias. Another category of counterfactual-based datasets evaluates bias by presenting the model with pairs of sentences containing different demographic terms and assessing their likelihood using a pseudo-log-likelihood metric. For instance, CrowS-Pairs (Nangia et al., 2020) provides 1,508 sentence pairs to assess stereotypes of historically marginalized social groups, addressing nine biases related to race, gender, sexual orientation, religion, age, nationality, disability, physical appearance, and socioeconomic status. Similarly, WinoQueer (Felkner et al., 2023), developed by the LGBTQ+ community, contains 45,540 sentence pairs designed to gauge anti-LGBTQ+ stereotypes, encompassing LGBTQ+ identity descriptors and counterfactual versions without LGBTQ+ identity markers.
@@ 9. Conclusion
LLMs have demonstrated remarkable success across various high-impact applications, transforming the way we interact with technology. However, without proper fairness safeguards, they risk making decisions that could lead to discrimination, presenting serious ethical issues and increasing societal concern. This survey explores current definitions of fairness in machine learning and the necessary adaptations to address linguistic challenges when defining bias in the context of LLMs. Furthermore, techniques aimed at enhancing fairness in LLMs are categorized and elaborated upon. Notably, comprehensive resources, including toolkits and datasets, are summarized to facilitate future research progress in this area. Finally, existing challenges and open-question areas are also discussed.
Acknowledgement
This work was supported in part by the National Science Foundation (NSF) under Grant No. 2245895.